# Sketch-Guided Constrained Decoding for Boosting Blackbox Large Language Models without Logit Access

### v1

@@ 4 Related work
[…]
Collaborative generation. In the realm of collaborative generation, Vernikos et al. (2023) and Welleck et al. (2023) explored training smaller language models to refine the outputs from larger models for enhanced quality. The skeleton-of-thought approach, proposed by Ning et al. (2023), generates an initial output skeleton and then concurrently develops each segment. Meanwhile, grammar prompting, as suggested by Wang et al. (2023), involves creating a meta-grammar to guide the output of LLMs in producing valid results.
In contrast to these methods, SGCD offers a versatile approach that is not confined to any specific API design. It is particularly aimed at enabling constraint application to the outputs of blackbox LLMs, addressing a distinct challenge in the field.
@@ 5 Conclusion
We propose sketch-guided constrained decoding (SGCD), a simple method for constrained decoding with blackbox LLMs that does not require access to next-token logits during generation. Our work is complementary to existing methods for constrained decoding and can be used in conjunction with them. Despite its simplicity, SGCD achieves strong performance on constrained generation tasks, outperforming vanilla LLMs (both blackbox and open-source) by a large margin.
@@ 6 Limitations
  • Limitation 1: The degeneration problem sometimes hampers SGCD. More investigation is needed to understand its causes and how to mitigate it.
  • Limitation 2: As LLMs keep getting better, the benefits of SGCD might diminish on some tasks. This is, however, not a limitation of SGCD per se, but rather of constrained decoding in general.
@@ Acknowledgments
We thank Grant Slatton for insightful discussions during the ideation phase. West’s lab is partly supported by grants from Swiss National Science Foundation (200021_185043, TMSGI2_211379), Swiss Data Science Center (P22_08), H2020 (952215), Microsoft Swiss Joint Research Center, and Google, and by generous gifts from Facebook, Google, and Microsoft.
@@ D.1 Task instruction
(b) Instruction for constrained generator
@@ D.3 IE datasets
The SynthIE-text dataset, synthesized by reverse prompting Text-Davinci-003 with triplets from Wikidata, stands out due to its substantial size, diverse content, and high-quality human ratings, as highlighted in Josifoski et al. (2023). This contrasts with prior datasets such as REBEL Huguet Cabot and Navigli (2021), whose annotation quality is low (Josifoski et al., 2022). However, a potential minor bias may exist towards GPT-4 and GPT-3.5, as SynthIE-text was generated from a model in their family, Text-Davinci-003. Despite this, we maintain that this does not compromise the validity of our method, given that our primary focus is on the comparative performance with and without the application of SGCD.
@@ D.4 Discussion of GPT-4 on cIE
An intriguing finding is that SGCD’s performance on the SynthIE-text dataset using GPT-4 (F1=45.6) is marginally lower than that achieved with few-shot prompting alone, without constrained decoding (F1=45.8). Our analysis suggests that the constrained generator occasionally struggles to adhere to the sketcher’s outline, resulting in fewer triplets than expected in the output. This observation is consistent with Wang et al. (2023)’s findings, where constrained decoding was noted to reduce the diversity of the generated samples.
@@
However, in real-world scenarios, text is typically more intricate, and grounding entities and relations in the KG is not as straightforward. Despite this, the overall performance enhancement provided by SGCD across various models remains noteworthy, averaging gains of up to 10.7% and 8.1% on Wiki-NRE and SynthIE-text, respectively.
@@ E.1 Task instruction
(b) Instruction for constrained generator
@@ E.2 Constraints and grammar
Grammar. The tree structure of the parse tree is usually captured by a context-free grammar. However, the context-free grammar is not sufficient to capture the completeness constraint, motivating the use of a more restrictive grammar. We use the same grammar as in Geng et al. (2023) as shown in Fig. 6.
@@
Figure 6: Grammar for constituency parsing used in constrained decoding.
@@
E.3 Example of constituency parsing correction
We provide an example to show the correction from our method in Figure 7. As shown in the figure, the parse tree of the input sequence is not valid because the number of closing brackets is 2 times less than the number of opening brackets and some tags are not consistent with the Penn Treebank format. The constrained generator generates a sequence of tokens that is consistent with the grammar and fixes the parse tree to be valid.
However, the fixed parse tree is not yet identical to the target parse tree, cf. blue nodes in Figure 7. This is expected because the constrained generator can only fix the parse tree to be consistent with the grammar but cannot guarantee that the parse tree is identical to the target parse tree.
Figure 7: Example to show the correction process of our method. The red nodes are the nodes that are corrected by constrained generator. The blue nodes are the nodes that different from the target parse tree( some of them are not corrected by constrained generator because they are not violating the grammar).
E.4 Validity adjusted metric
We use the validity adjusted metric as done in Hemmer et al. (2023) to measure the effectiveness of our method. It is defined as follows:
$\text{Metric}^{\text{v}}=\frac{1}{N}\sum_{i=1}^{N}\mathbbm{1}(\text{valid}(y_{i}))\cdot\text{Metric}(y_{i})$
$\text{valid}(y_{i})=\begin{cases}1&\text{if the parse tree is valid}\\ 0&\text{otherwise}\end{cases}$

### v3

@@ 4 Related work
Constrained decoding. Deutsch et al. (2019) introduced a general constrained decoding framework for text generation based on automata. Scholak et al. (2021); Poesia et al. (2022); Geng et al. (2023) implemented incremental parsing for domain-specific tasks such as SQL generation. Beurer-Kellner et al. (2023); Poesia et al. (2023) have proposed iterative approaches to constrained decoding using blackbox LLM APIs, albeit with potential limitations such as excessive API calls (thus increasing monetary cost), as detailed in Appendix C.
Collaborative generation. Vernikos et al. (2023) and Welleck et al. (2023) explored training smaller language models to refine the outputs from larger models for enhanced quality. The skeleton-of-thought method (Ning et al., 2023) generates an initial output skeleton and then concurrently develops each segment. Grammar prompting (Wang et al., 2023) creates a meta-grammar to guide the output of LLMs in producing valid results.
@@ 5 Conclusion
So far, constrained decoding has been limited to open-source models that provide access to their logits during generation. Overcoming this limitation, we propose sketch-guided constrained decoding (SketchGCD), a simple method for constrained decoding with blackbox LLMs that does not require access to next-token logits during generation. By using separate sketching and refinement phases, SketchGCD allows to benefit from the power of blackbox LLMs while still enforcing constraints. Our work is complementary to existing methods for constrained decoding and can be used in conjunction with them. Despite its simplicity, SketchGCD achieves strong performance on tasks exhibiting strong structural constraints, outperforming unconstrained generation by a large margin.
@@ 6 Limitations
The limitations of our method include the following. First, SketchGCD adds an overhead as it requires a constrained decoder to refine the sketches after the sketching phase. Second, as LLMs keep getting better, the benefits of SketchGCD might diminish on some tasks as the unconstrained model’s performance improves. Third, just as classical constrained decoding, SketchGCD can only enforce constraints at the structure level or the syntactic level, but not at the semantic level. The model can still generate semantically incorrect outputs. However, in many real-world applications, we have observed semantic errors to be less common than structural errors.
@@ Acknowledgments
We thank Grant Slatton for insightful discussions during the ideation phase. West’s lab is partly supported by grants from Swiss National Science Foundation (200021_185043, TMSGI2_211379), Swiss Data Science Center (P22_08), H2020 (952215), Microsoft Swiss Joint Research Center, and Google, and by generous gifts from Meta, Google, and Microsoft.
@@ D.1 Task instruction
(b) Instruction for constrained decoder
@@ D.3 IE datasets
The SynthIE-text dataset, synthesized by reverse prompting Text-Davinci-003 with triplets from Wikidata, stands out due to its substantial size, diverse content, and high-quality human ratings, as highlighted in Josifoski et al. (2023). This contrasts with prior datasets such as REBEL Huguet Cabot and Navigli (2021), whose annotation quality is low (Josifoski et al., 2022). However, a potential minor bias may exist towards GPT-4 and GPT-3.5-Turbo, as SynthIE-text was generated from a model in their family, Text-Davinci-003. Despite this, we maintain that this does not compromise the validity of our method, given that our primary focus is on the comparative performance with and without the application of SketchGCD.
@@ D.4 Discussion of GPT-4 on cIE
An intriguing finding is that SketchGCD’s performance on the SynthIE-text dataset using GPT-4 (F1=45.6) is marginally lower than that achieved with few-shot prompting alone, without constrained decoding (F1=45.8). Our analysis suggests that the constrained decoder occasionally struggles to adhere to the sketcher’s outline, resulting in fewer triplets than expected in the output. This observation is consistent with Wang et al. (2023)’s findings, where constrained decoding was noted to reduce the diversity of the generated samples.
@@
However, in real-world scenarios, text is typically more intricate, and grounding entities and relations in the KG is not as straightforward. Despite this, the overall performance enhancement provided by SketchGCD across various models remains noteworthy, averaging gains of up to 10.7% and 8.1% on Wiki-NRE and SynthIE-text, respectively.
@@ E.1 Task instruction
(b) Instruction for constrained decoder
@@ E.2 Constraints and grammar
Simple Context-Free Grammar. The tree structure of the parse tree is usually captured by a context-free grammar as shown in Table 8.
root ::= tree;
tree ::= node;
node ::= clause | phrase | word;
clause ::= spaced_open_parenthesis, space,
            clause_tag, function_tag*,
            index?, node*,
           spaced_close_parenthesis;
phrase ::= spaced_open_parenthesis, space,
            phrase_tag, function_tag*,
            index?, node*,
            spaced_close_parenthesis;
word ::= spaced_open_parenthesis, space,
        word_tag, space, actual_word,
        spaced_close_parenthesis;
clause_tag ::= "S" | ...  | "SQ";
phrase_tag ::= "ADJP" | ...| "WHADVP";
word_tag ::= "CC" |...|"WRB";
function_tag ::= "-ADV" |... | "-TTL";
actual_word ::= "xxx";
index ::= "-", [1-9], {0-9};
spaced_open_parenthesis ::= space, "(";
spaced_close_parenthesis ::= space, ")";
space ::= " ";
Table 8: Lite Context-Free Grammar for constituency parsing.
Sophisticated Regular Grammar. However, the context-free grammar is not sufficient to capture the completeness constraint, motivating the use of a more restrictive grammar. Geng et al. (2023) proposed a sophisticated regular grammar to enforce the constraints of completeness, balanced brackets, and label consistency as shown in Fig. 6.
@@
Figure 6: Sophisticated Regular Grammar for constituency parsing.
@@
Method Bracket-Prec Bracket-Recall Bracket-F1 Tag Accuracy Valid Tag Valid Tree
GPT-4 76.6 67.7 71.9 95.4 93.6 86.0
       + Lite Context-Free Grammar 75.8 67.8 71.5 95.3 100 92.5
       + Sophisticated Regular Grammar 69.3 63.1 66.1 98.5 100 100
GPT-3.5-Turbo 68.2 55.5 61.2 93.1 91.7 76.9
       + Lite Context-Free Grammar 68.7 56.6 62.1 92.6 100 81.5
       + Sophisticated Regular Grammar 61.2 49.4 54.7 96.0 100 100
Claude 73.1 63.1 67.7 94.5 95.1 62.6
       + Lite Context-Free Grammar 71.6 62.9 66.9 93.4 100 68.7
       + Sophisticated Regular Grammar 52.1 45.4 48.5 75.9 100 99.2
Claude-instant 71.3 59.1 64.7 89.6 91.7 56.6
       + Lite Context-Free Grammar 66.6 57.4 61.6 87.9 100 67.8
       + Sophisticated Regular Grammar 59.6 49.2 53.9 84.9 100 99.5
Table 9: Constituency parsing with two different grammar constraints, measured in terms of bracketing recall, precision, F1-score, and tag accuracy (with bootstrapped 95% confidence intervals) †Only subset of samples whose ground-truth parse trees are shorter than 128 tokens(LLaMAtokenizer) are considered, which accounts for shortest 25% of the samples.